OpenStudy (anonymous):

6/(x^2+7x)-4/(x^2-49)

Parth (parthkohli):

Can you factorize the denominators?

OpenStudy (anonymous):

\[\frac{ 6 }{ x^2+7x}-\frac{ 4}{ x^2-49}\]

OpenStudy (anonymous):

i need to find the differnece

OpenStudy (anonymous):

\[\frac{ 2(x-21) }{ x(x+7)(x-7)}\] is what i got

OpenStudy (anonymous):

\[\frac{ 2x }{5x^2-18x+16 }-\frac{ 8x }{ 7x^2-18x+8}\]